How many years will it take $50,000 to grow to $500,000 in an account that pays 4.5% compounded daily?
Answer #1

We use the formula:
A=P(1+r/365)^365n
where
A=future value
P=present value
r=rate of interest
n=time period.

500,000=50,000*(1+0.045/365)^365n

(500,000/50,000)=(1.000123288)^365n

Taking log on both sides;

log 10=365n*log (1.000123288)

n=1/365[log 10/log (1.000123288)]

=51.17 years(Approx).